Journalist Report – February 18th

Journalist report, 02/17/18 and 02/18/18: last moments on Earth

Louis had been the last to join us at Grand Junction, picked up by Gabriel, glad of his time of arrival. The mission was finally on its way! After a final brunch at the hotel, Victoria made a stock of bananas (which would be over even before we could see the Hab), it was time to leave civilization! The target: Hanksville, in the deepest paths of the Utah desert.

A journey not so comfortable at last, especially for Laurent, who carried the large MegaARES instrument that Gabriel left him. Nonetheless the road met its end, and we were getting deeper in the desert, towards the Hab. A few turns between giant red rocks later, we finally were able to see the station, among its little brothers, the laboratory, the GreenHab and the observatory. To welcome us, no previous crew was present contrary to the usual, but new stairs were being done, which will reduce falling risks and avoid injuries. It was then time for Shannon’s first briefing, completed with some advices from Attila the Peruvian, yet hard to understand because of Jack the dog, which seems to be into his teenager crisis.

Following a small cleaning session imposed by Victoria, it was now time for all of us to choose our rooms. Well, except for Gabriel, who inherited of the small bed under the celling, in order to take care of the hydraulic pump. A small rover/ATV trip later and we were taking our first meal in the station! A meal which finally revealed to be quite tasty: lyophilised ingredients are not so terrible. However, the crew started to feel really tired and everyone went to sleep to prepare the last pre-mission day!

A sleeping night which was short for Benoît and Louis, who had to bring back the location car to Grand Junction: it is an early morning for them! A few hours later, the remaining members of the crew tasted their first Martian breakfast with powdered orange juice and pancakes … powdered pancakes. The morning is used to put in place experiments which will be completed during the whole 3 weeks, and to finish some secondary tasks, like patch sewing, realized by Victoria, myself and Benoît for diverse results …

As both drivers came back from their journey and everyone filled their stomach, it was the time to take crew pictures! Despite some difficulties (one dog which repeatedly entered the field of view, and a tragic fall of the camera on the ground), photos were nonetheless satisfying. A last rover/ATV trip was organized, which proved to be quite hard due to continuous strong blowing winds and small dust storms that came with it, leading some members of the crew (especially Gabriel) to steal the look of a grandmother with his bandana.

The fatidic hour was near: while Shannon gave us her last advices, we prepared ourselves to the closing of the Hab door. Here begins 3 weeks without seeing the Sun without the helmet. As Shannon wished us good luck and left us, Gabriel finally grabbed the door handle to close the main door. After a few videos and pictures were taken by Jeremy, it was at last time to start the simulation.

6.05pm: the Hab door is closed, here begins the mission

Alexandre Martin, Crew Journalist MDRS 189.

Commander Report – February 18th

Crew 189 Commander Report 18Feb2018

Sol 0
Today began early for Benoit and I: because no crew was in the MDRS
before us to bring it back, we had to bring the rental car back to
Grand Junction. During this time, the rest of us had a bit of free
time to rest, then was trained by Dr Rupert on the Hab habits and map
use. We then had lunch together, and started the afternoon by our crew
official photo, in front of the hab. We then tried go on the ATVs but
quickly ended up in a little sand storm, so that we didn’t go very
far.
After a last briefing more centered on simulation philosophy, came the
time to officially close the airlock. Even if it was less magical than
last year, it was still very symbolic: we are now by ourselves, and
for 3 weeks.
Everything is now slowly getting in order: everybody finds his role,
we wrote our first reports today, tested various equipment and look
forward to our first EVA, if the weather permits it! We want both to
train our new crewmembers about EVA specifications, and to find an
acceptable spot to display our experiments. As these will need to
remain on the field as long as possible, we want to make it a priority
for the beginning of the mission.
Everything is going pretty fast at the moment, but I think tomorrow
will be our last day of rush, and that after it, everybody will be
properly settled, and ready for the real business. Anyway, I see
excitement in all of us, and it is a good thing: everybody seems happy
to be here and motivated. I look forward to seeing what this crew is
going to accomplish!

Louis Mangin

Operations Report – February 18th

Crew 189 Operations Report 18Feb2018

SOL: 0

Name of person filing report: Gabriel PAYEN

Non-nominal systems: Pipe leak in low deck, water heater, two thermometers in the Greenhab.

Notes on non-nominal systems: –

Generator (hours run): 14 hours during the night (7 PM to 9 AM)

Solar— SOC % (Before generator is run at night) : 91%

Diesel – ?

Propane – 46.2 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 7 gallons (estimation based the total height and capacity of the container)

Water (trailer) – 0 gallons.

Water (static) – 400 gallons (estimation based the total height and capacity of the tank)

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter (loft) : 0.44, measurement at 7 PM

Toilet tank emptied: No

ATV’s Used: All ATVs used except the one which doesn’t work (I forgot to check its number)

Oil Added : No

ATV Fuel Used: Gals

# Hours the ATVs were Used today: 20min

Notes on ATVs: –

Deimos rover used: Yes

Hours: 108.7

Beginning charge: ? (not checked)

Ending charge: 98%

Currently charging: No

Spirit rover used: No

Hours: 20.2

Beginning charge: –

Ending charge: –

Currently charging: Yes (Atilia may have unplugged Curiosity to plug in Spirit)

Opportunity rover used: Yes

Hours: 15.3

Beginning charge: ? (not checked)

Ending charge: 93% (100% after being charged)

Currently charging: No

Curiosity rover used: No

Hours: 12.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

HabCar used and why, where : Not used

General notes and comments: Last checks today with Shannon, some things to fix (see Non-nominal systems). Jérémy tried to run his aerosol counter device (LOAC) in the Hab but it doesn’t work. It turns on but doesn’t begin counting. The crew will try to fix it tomorrow.

Summary of internet: –

Summary of suits and radios: All air pipes had been tested and only one doesn’t work (backpack 3 – left one). All radios had been tested and work

Summary of Hab operations: –

Summary of GreenHab operations: –

Summary of ScienceDome operations: –

Summary of RAMM operations: –

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support:

-Thermometers

The GreenHab officer would need a new thermometer.

-Pipe leaks

A bucket is beneath the pipe and receives a liquid which seems to be bad smelling water. Tape would stop the flow, but it only seems to be a temporary solution. Tomorrow I’ll try to find where the leak comes from.

-Water heater

Shannon seems to have fixed it before the very beginning of the SIM. We’ll see it when a crew member will take the first shower of the mission.

-Others:

Some checks have been forgotten (beginning charge of the rovers, numbers of ATVs used…). Sorry for that, I’ll try not to do theses mistakes anymore.

GreenHab Report – February 18th

GreenHab Report

AUCLAIR Jérémy

18 feb 2018

Environmental control:

Ambient with fan (high winds today)

Heating

Shade cloth on

Average temperatures:

Both thermometers are out of service

I estimated the temperature at 35°C around mid-day

Hours of supplemental light: from 8pm to 11pm

Changes to crops: none

Daily water usage for crops: around 15 gallons

Time(s) of watering for crops: 11am and 6pm

Narrative:

Functionality: The GreenHab heaters are working fine. Both thermometers are out of service, but around mid-day I estimated the temperature at approximately 35 °C, I turned on the fan for half an hour to keep the GreenHab cool. Because of high wind I left the door closed. Humidity was low in the morning (11am – 21 %) and rose up to 28 % during the day. I watered the existing seedlings twice (11am and 6pm) today and they are growing well.

Status: The existing seedlings in the green Hab are continuing to grow. There are a lot of tomato plants around the Hab, and some spinach, lettuce, beans, carrots and onion growing in the small pots. There are some other plants like a variety of spices and a few cucumbers. In the large rectangular pots there are a lot of salads (romaine mainly). I barely got familiar with the GreenHab today.

Planned activity: I will try to do the inventory stock of the greenhab tomorrow if I have time. I will look at the different seeds there are to choose some to plant them.

Support/supplies needed: one or two thermometers

Sol Summary – February 18th

Crew 189 Sol Summary Report 18 Feb 2018

Sol 0

Summary Title: Getting Ready

Author: Louis Mangin

Mission Status: All nominal

Sol Activity Summary: Last briefings, beginning of the simulation. We
started by a round trip to Grand Junction to get back the rental car
while the rest of the crew was settling in the hab, learned about the
surroundings and station management with Dr Rupert. Then we learned
about driving, spacesuits, and had a briefing about simulation with Dr
Rupert. We then closed the airlock at 5:50 PM before sunset.

Look Ahead Plan: Because of the wind planned for tomorrow, we changed
of plans, and will only go out for an EVA in the Hab surroundings to
select a location to deploy our air sensors, MegaARES and LOAC.

Anomalies in work: Water leak in the Hab Lower deck, broken fans in
spacesuits, water heater badly calibrated. Two greenhab thermometers
probably broken. Voltmeter not working.

Weather: Very windy, hard to see and to travel without space suits.

Crew Physical Status: Fine, already a bit tired from travel.

EVA: None

Reports to be file:
Commander
Journalist
HSO checklist
Engineering
Greenhab Officer

Support Requested: One or two thermometers.